A homeowner wishes to insulate her attic with fibergrlass insulation to conserve energy. The insulation comes in 40-cm wide rolls that are cut to fit between the rafters in the attic. if the roof is 6m from peark to eave and the attic space is 2 m high atthe peak, how long does each of the pieces of insulation need to be? round to the nearest tenth.
Answer by checkley71(8403)

what you have is a right triangle with the one side of 6 meters & one side of 2 meters.
using the pythagoreon theorum a^2+b^2=c^2 you get:
2^2+6^2=c^2
4+36=c^2
40=c^2
c=sqrt40
c=6.3 meters is the length of the insulation.